OpenStudy (anonymous):

what is the force required to move a load of 50 kg through a distance of 1000cm in 5 seconds

OpenStudy (akashdeepdeb):

We know that force = mass * acceleration We have mass but we need to find the acceleration needed to be produced by the force to move the object over a distance of 10m in 5s Take the initial velocity of the load = 0 m/s Distance to be covered = 1000cm = 10m Time = 5s Acceleration = ? Equation of motion to be used = \[s = ut + \frac{at^2}{2}\] Now plug in the values a = 4/5 = 0.8 \(m/s^2\) Now we have the acceleration and the mass. Thus, you can now find the force = ma required to push this 50 kg load over a distance of 10m in 5s. Understood? :)

OpenStudy (anonymous):

thank u

OpenStudy (akashdeepdeb):

:)